Find the equation of the line through the point (5,0) and perpendicular to the line y=(4x/3)+5. Express the equation of the line in slope-intercept form, y=mx+b.
So I am assuming that I need to put the points (5,0) into the y=mx+b so I did that. 0=(4/3)(5/1)+b and solved.
0=(20/3)+b so then I solved again so is b=-(20/3)?
I don't know what part of this I am doing wrong but I want to make sure I understand the content. Any help is greatly appreciated :)

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find the equation of the line through the point (5,0) and perpendicular to the line y=(4x/3)+5. Express the equation of the line in slope-intercept form, y=mx+b.
So I am assuming that I need to put the points (5,0) into the y=mx+b so I did that. 0=(4/3)(5/1)+b and solved.
0=(20/3)+b so then I solved again so is b=-(20/3)?
I don't know what part of this I am doing wrong but I want to make sure I understand the content. Any help is greatly appreciated :)
-------------------------------
The slope of lines perpendicular to the given line is the negative inverse, -3/4.
Then, y-y1 = m*(x-x1) where (x1,y1) is the given point.
y - 0 = (-3/4)*(x - 5)
y = (-3/4)x + 15/4

Answer by ikleyn(52779)

.
At this web-site,  there is a group of lessons
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
that are closely related to your problem.

They cover the entire subject with all inter-related problems in one way.

You have a happy opportunity to learn the subject from these lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.